OpenStudy (anonymous):

please help!!!! given the formula, solve for the indicated variable A=P(1+nr) for r

OpenStudy (lgbasallote):

divide both sides by P A/P = 1 + nr subtract 1 from both sides A/ P - 1 = nr divide both sides by n A/Pn - 1/n = r simplify... (A - P)/n = r

OpenStudy (lgbasallote):

the simplification is (A-P)/(Pn) = r sorry

OpenStudy (anonymous):

thank you so much